### v4

@@ Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully. We address this open problem by framing it as an “Embed-Search-Align” task. In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. DNA-ESA is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. DNA-ESA far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
@@ 1 Introduction
Sequence alignment is a central problem in the analysis of sequence data. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long. Moreover, most experiments will generate millions of reads. Therefore, the resulting task of aligning reads to genomes is computationally expensive, and many decades of work have led to optimized approaches that can perform these tasks with great efficiency. However, many limitations remain with existing genome alignment methods. One limitation is that a reference genome usually is an approximation of the sequence from a single individual, and does not capture the variability of genomes across populations. Another, is that sequence alignment algorithms may not correctly model sequencing errors, and thus may introduce mistakes when trying to score the best match for a sequence against a genome. We asked whether some of these limitations could potentially be overcome if we utilize a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
@@
We establish a foundation model foundation tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case). Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as genomes) and natural languages with large vocabulary. In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. Existing foundation Transformer models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. There are, however, tasks such as Sequence Alignment where the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective. This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LVLs to a continuous representation space.
@@
Figure 2: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA. In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings. Vertical marker lines (||||) refer to fragments, “–” markers correspond to reads. We observe that the consecutive fragments belonging to the same nucleotide (subfigure (a)) or gene (subfigure (b)) sequence constitute an order-preserving 1D manifold. Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
@@
In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020, ), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
1.1 Transformer models: Written Language to DNA Sequence Alignment
@@
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences. Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, subsets of genes.
@@
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function. Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
@@
2 Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space. (B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/. This provides efficient access to the entire reference sequence for each read alignment. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 Task Definition
@@
$v^{*}\approx\min_{\mathcal{F}_{i}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (4)
@@
    There are enough fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e. $\cup\mathcal{F}_{j}=\mathcal{R}$;
@@
    $K$ is significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$ then this amounts to scanning the whole reference.
@@
2.2 Designing effective sequence representations
Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
2.3 Self-supervision and contrastive loss
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work simclr . Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
2.4 Encoder implementation
DNA-ESA uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
2.5 Search and retrieval
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
Retrieval: Given a read $r$, we project its corresponding DNA-ESA representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
@@
Fine-Alignment: A standard SW distance library (biopython, ) is used to solve (4), which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
@@
3 Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 1. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(nucleotide, ), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(ji_dnabert_2021, ), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(nguyen_hyenadna_2023, ). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 1. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with DNA-ESA in Table 1.
DNA-ESA convergence plots are presented in Figs. 6(a) and  6(b). Model checkpoints are available at OSF. In Fig. 2, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (umap, ). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (6) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. The codebase is linked.
4 Sequence alignment of ART-simulated reads
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.

### v5

@@ Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison. We bridge this gap by framing the Sequence Alignment task for Transformer models as an “Embed-Search-Align” task. In this framework, a novel Reference-Free DNA Embedding RDE model generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, we introduce: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. RDE is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines on ESA such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
@@
Sequence alignment is a central problem in the analysis of sequence data. It is used in various genomic analyses, including variant calling, transcriptomics, and epigenomics. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long. Moreover, most experiments will generate millions of reads. The resulting task of aligning short reads to large-size genomes is computationally challenging –akin to finding a pin in a haystack– and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
Since genomes are sequences where the alphabet consists of only a few symbols ($\{A,T,G,C\}$) they can be considered as Limited Alphabet Languages (LAL). We ask whether one can design a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
@@
Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Alphabet Languages (LAL) (such as genomes) and natural languages with large vocabulary. In LALs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective. This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LALs to a continuous representation space.
We establish a foundation model foundation Reference-free DNA Embedding (RDE) – tailored for embedding DNA sequences. The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths – for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps – then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors. Another requirement of such RDE models is that they should generate embeddings that are reference free, i.e. once trained, it creates semantically-aware embedding (i.e. sequences that are close in edit distance are mapped close to each other) of any given genome subsequences irrespective of the reference genome from which it is sampled. Given such a model, alignment of reads can be achieved by a near-neighbor search in the embedding space: only the fragments of genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long sequence is reduced to a local search in a vector space.
@@
Figure 2: Illustrating RDE’s Preservation of Sequence Locality in Embedding Space: The reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using embeddings produced by RDE. In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings. Vertical marker lines (||||) refer to fragments, “–” markers correspond to reads. We observe that the consecutive fragments belonging to the same nucleotide (subfigure (a)) or gene (subfigure (b)) sequence constitute an order-preserving 1D manifold. Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
@@
In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
1.1 1.1 Transformer models: Written Language to DNA Sequence Alignment
@@
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Traditional DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
@@
Recognizing such similarities, within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function.
These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. A natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
@@
2 2. Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder Reference-Free DNA Embedding (RDE), trained through self-supervision, to embed DNA subsequences to vectors. (B) Next, we formulate a framework, Embed-Search-Align (ESA), that maps reads to reference genome sequence. We leverage a specialized data structure, termed a DNA vector store, that enables efficient storage and search in an embedding space [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/: the entire reference genome is sharded into equal-length overlapping fragments whose embeddings are then uploaded to the vector store. For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithimically in the number of fragments. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 2.1 RDE: Designing effective sequence representations
Subscript $j$ serves to indicate that the read $r_{j}$ is correctly aligned to a fragment reference, $\mathcal{F}_{j}$, (that we call a positive {read, fragment} pair). Given another arbitrary fragment $\mathcal{F}_{i}$ to which $r_{j}$ has poor alignment, $\{r_{j},\mathcal{F}_{i}\}$ is a negative {read, fragment} pair. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
2.2 2.2 RDE: Self-supervision and contrastive loss
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work chen_simple_2020 . Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
2.3 2.3 RDE: Encoder implementation
RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
2.4 2.4 ESA: Task Definition
@@
$v^{*}\approx\min_{\mathcal{F}_{j}\in\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}}\texttt{SA}\left(r,\mathcal{F}_{j}\right).$ (6)
@@
    Since a read can originate from any position along the reference, there must be sufficient fragments $\mathcal{F}_{i}$ to cover $\mathcal{R}$, i.e. $\cup\mathcal{F}_{j}=\mathcal{R}$. From within this set of fragments, we hope to retrieve a subset of fragments containing the optimal fragment for the read;
@@
    While retrieving a subset of fragments of size $K$, we require $K$ to be significantly smaller than $\frac{N}{Q}$. If $\frac{N}{Q}$, then this amounts to scanning the whole reference.
@@
2.5 2.5 ESA: Search and retrieval
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
Retrieval: Given a read $r$, we project its corresponding RDE representation into the vector store and retrieve the approximate nearest-$K$ set of reference fragment vectors and the corresponding fragment metadata $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$.
@@
Fine-Alignment: A standard SW score library (biopython, ) is used to solve (6), which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
@@
3 3. Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 1. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(nucleotide, ), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(ji_dnabert_2021, ), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(nguyen_hyenadna_2023, ). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads (reads without noise in comparison to the reference) of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 1. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with RDE on ESA in Table 1.
RDE convergence plots are presented in the SM Sec. A. Model checkpoints are available in OSF osf . In Fig. 2, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (umap, ). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (4) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The RDE representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. The codebase is linked.
4 4. Sequence alignment of ART-simulated reads
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, RDE and the two best baselines – DB2, mean and HD, max – are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.